OpenStudy (anonymous):

write a mfile with matlab which takes a and n from user and gives the n term s of taylor series of sin(x) around x=a

OpenStudy (anonymous):

i have difficulties with math part @experimentX

OpenStudy (anonymous):

we must print something like this\[a_0+a_1(x-a)+a_2(x-a)^2+...\]

OpenStudy (anonymous):

man i have nothing :(

OpenStudy (experimentx):

Woops ... !! probably write a function do do that.

OpenStudy (experimentx):

syms x diff(sin(x), 'n times differentiate', 'around a');

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

man would u mind to join us :)

OpenStudy (experimentx):

Woops!! lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

nth derivative of sin(x)\[\sin(\frac{n\pi}{2}+x)\]

OpenStudy (anonymous):

wrong !!!

OpenStudy (experimentx):

Woops ... i need to differentiate it and substitute it.

OpenStudy (experimentx):

clc syms x center = input('Enter the center of expansion: '); for i=1:10 disp(subs(diff(sin(x), i), x, center)); end

OpenStudy (anonymous):

exper can we show that like what i mentioned above?

OpenStudy (anonymous):

for example with center 3 -0.9900-0.1411(x-3)+0.9900(x-3)^2+...

OpenStudy (experimentx):

hold on ... I'm working.

OpenStudy (experimentx):

I don't like Matlab's string handing !!

OpenStudy (anonymous):

me too

OpenStudy (phi):

This does not do the input. But it does give an idea a=pi/3, n=4; str2=''; for ii=0:(n-1) switch mod(ii,4) case 0, ff= sin(a); case 1, ff= cos(a); case 2, ff= -sin(a); case 3, ff= -cos(a); end; if (ff<0) sg= '-' else sg= '+'; end; if ii>0 str= sprintf('%2s %7.5f(x-%2.4f)^%d',sg, abs(ff)/factorial(ii),a,ii); else str= sprintf('%7.5f',ff); end; str2= [str2 str]; end; str2

OpenStudy (phi):

its output str2 = 0.86603 + 0.50000(x-1.0472)^1 - 0.43301(x-1.0472)^2 - 0.08333(x-1.0472)^3

OpenStudy (experimentx):

I really hate matlab for string manipulation clc syms x center = input('Enter the center of expansion: '); for i=1:3 disp( [subs(diff(sin(x), i), x, center), strcat('(x-', num2str(center), ')^', num2str(i))]); end

OpenStudy (experimentx):

I got this nasty code ... looks pretty close clc syms x center = input('Enter the center of expansion: '); to_out = ''; for i=1:10 to_out = strcat(to_out, '+' , strcat(char(subs(diff(sin(x), i), x, center)), strcat('(x-', num2str(center), ')^', num2str(i)))); end disp(to_out);

OpenStudy (experimentx):

the out put is ...for r=4 ... maybe tweak a little. +cos(3)(x-3)^1+-sin(3)(x-3)^2+-cos(3)(x-3)^3+sin(3)(x-3)^4

OpenStudy (anonymous):

thank u guys